# Multi-Robot Rendezvous in Unknown Environment with Limited Communication

### v1

@@ I Introduction
Multi-robot systems have been receiving increasing attention in recent years [1, 2]. The cooperative motion in multi-robot systems has been widely studied, for example, consensus [3], exploration [4], and rendezvous [5]. Among the above-mentioned cooperative motion, rendezvous is one of the most important topics, which aims at gathering all robots at a specific location in the environment at a given moment [6]. This behavior is common in the real world, for example, a few friends want to meet with each other in a shopping mall or a rescue helicopter is looking for lost hikers in the desert [7].
@@
One of the simplest rendezvous problems involves two robots moving along a straight line, which was first introduced in [16]. The solutions to this problem can be divided into two types: symmetric and asymmetric. The symmetric version requires the robots to execute the same search strategy. Symmetric rendezvous strategy for two robots switching among three fixed locations was proved to be optimal in [15]. Besides, it was extended to a continuous scenario in [8]. The asymmetric version was studied in [17, 18]. Rendezvous for $n>2$ robots on a line was studied in [7] and was extended to 2D scenario in [14, 13]. In these researches, robots rely on stochastic strategies for rendezvous. However, in scenarios with a large number of robots or large-scale environments, the random searching strategy makes rendezvous very challenging or even infeasible.
@@
The condition of communication was introduced in [19, 5, 20, 9, 21], where the environment is usually assumed to be known and robots can locate themselves in the environment. Moreover, RPs between robots are available under this assumption. This kind of rendezvous problem has been receiving significant attention in control theory. However, rendezvous in unknown environments is more common and valuable, which is rarely studied under the assumption of limited communication. As far as we know, this is the first work to perform rendezvous under this assumption.
@@
Recently, multi-robot exploration with unknown initial RP has been studied [22, 4, 23]. In the process, robots determine their RPs and merge the constructed map. In theory, after the exploration and map construction, it is possible to use this full map to achieve rendezvous. However, simply applying multi-robot exploration algorithms to achieve rendezvous is inefficient. Firstly, the exploration aims to construct a complete and detailed map of the environment, while for rendezvous, exploring the environment as minimally as possible is preferred. Actually, an incomplete exploration is needed for rendezvous to improve the efficiency. Secondly, what kinds of information should be transmitted between robots has also not been studied to effectively achieve RP estimation and rendezvous point selection, which will lead to high communication cost.
@@
Thus, we propose a strategy called partitioned and incomplete exploration for rendezvous (PIER). We focus on how to effectively utilize the limited communication bandwidth (we assume communication range is not limited). This problem involves two aspects specifically: first, how to achieve RP estimation, and second, how to transmit environmental structural information to enable the selection of rendezvous points. In our previous paper, we proposed a lightweight Feature-based Hybrid Topological Map (FHT-Map) [24]. There are two types of nodes: main nodes and support nodes. Main nodes store features extracted from the environment and can facilitate RP estimation. Environmental structural information is stored in support nodes with very low storage requirements, which can be used for rendezvous point selection. By sharing lightweight FHT-Maps among robots, we can achieve rendezvous with low bandwidth requirement.
@@
Thus, we divide the rendezvous problem into two steps. Firstly, each robot performs rendezvous based environment exploration using PIER and constructs its FHT-Map. During this process, robots share FHT-Maps among each other and detect similar main nodes for RP estimation, and the FHT-Maps can be merged one by one. Secondly, once the RPs between all robots are determined, the optimal point for rendezvous can be selected on the merged FHT-Map. Our main contributions are:
@@
  • PIER is proposed for solving incomplete exploration task to achieve rendezvous.
@@ II Preliminaries
In this paper, we consider the rendezvous problem in a finite 2D space $S\subset\mathbb{R}^{2}$, where the size of the environment is unknown initially. This assumption is easy to satisfy [25], for example, in rendezvous problems for humans within a shopping mall or a campus. We define the k-th robot in the multi-robot system as $\mathcal{R}^{k}$ and the total number of robots is $n$. $\mathbf{P}^{k}=[\mathbf{p}^{k},\mathbf{q}^{k}]\in\mathbf{SE}(2)$ describes the pose of $\mathcal{R}^{k}$ in its own map frame, where $\mathbf{p}^{k}\in\mathbb{R}^{2}$ and $\mathbf{q}^{k}\in\mathbf{SO}(2)$ denote the position and orientation of $\mathcal{R}^{k}$. Besides, a distributed communication framework is used, where all communication occurs between two robots without a central server. An undirected communication graph $\mathcal{G}_{c}=(\mathcal{V}_{c},\mathcal{E}_{c})$ is used to describe the communication topology between $n$ robots. The node set is $\mathcal{V}_{c}=\{\mathcal{R}_{1},\cdots,\mathcal{R}_{n}\}$, the edge set is $\mathcal{E}_{c}\subseteq\mathcal{V}_{c}\times\mathcal{V}_{c}$. $\mathcal{N}_{c}^{k}$ defines the set of neighboring robots on $\mathcal{G}_{c}$ for $\mathcal{R}^{k}$.
@@
Different data can be used for RP estimation, for example, images [26] and point clouds [27]. Using raw data from the sensors for RP estimation provides more accurate results but imposes a significant burden on communication. In this paper, we employ lightweight Place Recognition (PR) descriptors [28, 29, 30] for rough RP estimation. Only when two PR descriptors are matched, 2D laser scans of these two points will be transferred for a more accurate RP estimation. An undirected RP graph $\mathcal{G}_{r}=(\mathcal{V}_{r},\mathcal{E}_{r})$ is used to describe the already estimated RPs between $n$ robots. The node set in $\mathcal{G}_{r}$ is $\mathcal{V}_{r}=\{\mathcal{R}^{1},\cdots,\mathcal{R}^{n}\}$, the edge set is $\mathcal{E}_{r}\subseteq\mathcal{V}_{r}\times\mathcal{V}_{r}$. Initially, since there are no estimated RPs between any two robots, $\mathcal{E}_{r}$ is an empty set. When the RP estimation is obtained between robots $m$ and $n$, the edge $(\mathcal{R}^{m},\mathcal{R}^{n})$ is added to $\mathcal{E}_{r}$. The details for RP estimation will be further described in Section III-B.
@@ III PIER Formulation
Robot exploration for a complete reconstruction of the environment has been widely investigated in recent years [23, 33, 4, 34]. In [23], Next Best View (NBV) is used for multi-robot exploration. NBV-based methods employ a greedy strategy, enabling rapid acquisition of environmental features but potentially resulting in poor cooperation between robots. In [33, 4], multi-robot exploration is modeled as a multiple traveling salesman problem (m-TSP), which optimizes the movements based on current frontiers. However, m-TSP-based methods are not efficient for the unknown space exploration. In [34], the Voronoi diagram is used for the partition of the space, and the collaborative behavior is enhanced during exploration. Compared with a detailed reconstruction, rendezvous based environment exploration aims at acquiring environmental features for RP estimation. Therefore, utilizing a greedy strategy to acquire features and considering cooperation between robots will be time-efficient for rendezvous. So, partitioned and incomplete exploration for rendezvous (PIER) is proposed in this paper. In PIER, we combine NBV-based and partitioned strategies to achieve rapid cooperative exploration by multiple robots for rendezvous under the condition of unknown initial RPs. In the following section, we will be describe the NBV-based exploration for single robot firstly. Then, the detailed methods for RP estimation and cooperative exploration will be presented.
@@ III-A Single Robot NBV Exploration
The NBV-based method employs a greedy strategy, which can achieve a faster environment coverage for incomplete exploration tasks. During this process, the environment $S$ can be divided into two parts: known space $S_{\text{known}}$ and unknown space $S_{\text{unknown}}$. Furthermore, it can also be classified into two types: free space $S_{\text{free}}$ and occupied space $S_{\text{occ}}$. In PIER, the k-th robot $\mathcal{R}^{k}$ is assigned with a portion of the area $S^{k}$ to explore. The combination of each space forms the environment $S^{1}\cup S^{2}\cdots\cup S^{n}=S$.
@@ III-B RP Estimation between Robots
To achieve rendezvous, each robot first performs NBV-based strategy to explore the unknown environment. During this process, each robot constructs an FHT-Map [24] which can be used for RP estimation. In the FHT-Map, there are two types of nodes: main nodes and support nodes. Main nodes store the visual features $\boldsymbol{\phi}^{k,i}\in\mathbb{R}^{512}$ and geometric features $\boldsymbol{\psi}^{k,i}\in\mathbb{R}^{360}$ of the location, which can be used for RP estimation. When $\mathcal{R}^{k}$ passes a main node of other robots, it can obtain the RP between $\mathcal{R}^{k}$ and that robot. So, it is crucial to select appropriate positions of main nodes to accelerate RP estimation and rendezvous. In this section, we will describe the details of using FHT-Maps for RP estimation.
@@ III-B1 Selection of Main Nodes
Secondly, since each robot performs incomplete exploration, the probability of each position in the environment being visited is different. An intuitive scenario is that robots are more likely to pass through crossroads in the environment, which is shown in Fig. 2. Therefore, selecting main nodes at these positions can facilitate RP estimation. We refer to the probability of a location $\textbf{p}^{k}(t)$ being visited as “heat value” $\mathcal{H}(\textbf{p}^{k}(t))\in\mathbb{R}$. We use a neural network to fit the heat value for each location. This network takes the laser scan $\boldsymbol{\psi}(\textbf{p}^{k}(t))$ as input and outputs the fitted $\mathcal{H}(\textbf{p}^{k}(t))$. Heat values of different locations are labeled manually based on Explore-Bench data set [35]. Besides, the backbone of this network is selected as the Pointnet [36].
@@
$\ \boldsymbol{\phi}(\textbf{p}^{k}(\tau))^{T}\boldsymbol{\phi}^{k,i}<th_{\text{sim}},\ \forall v_{f}^{k,i}\in\mathcal{V}_{f,\text{main}}^{k},$
where $th_{\text{sim}}$ is the similarity threshold between two locations [37]. $\gamma_{1}$ and $\gamma_{2}$ are two parameters which control the number of potential main nodes, which is the same as [24].
In Eq. 2, the feasible set of $\textbf{p}^{k}(\tau)$ is determined firstly. As $\mathcal{R}^{k}$ explores the environment, it gradually moves towards unknown areas, leading to a reduction in main nodes density $\mathcal{D}(\textbf{p}^{k}(t))$. Besides, when $\mathcal{D}(\textbf{p}^{k}(t))$ is smaller than $\gamma_{1}$, the robot will store all features it captures at a fixed rate (30 Hz). These features are called potential main nodes. When $\mathcal{D}(\textbf{p}^{k}(t))$ falls below $\gamma_{2}$, the single main node selection is performed. Firstly, potential main nodes are defined as those who satisfy $\gamma_{2}<\mathcal{D}(\textbf{p}^{k}(\tau))<\gamma_{1}$. Then, we do not expect that the selected main node has high similarity with the existing main nodes in the FHT-Map, which will lead to wrong estimations for RP. Potential main nodes are filtered using $\boldsymbol{\phi}(\textbf{p}^{k}(\tau))^{T}\boldsymbol{\phi}^{k,i}<th_{\text{sim}}$. Finally, a new main node will be chosen among the remaining potential main nodes.
Weighted sum method can be used to solve Eq. 2. The object function can be reshaped into
$\min_{\tau}\ \omega_{1}\mathcal{D}(\textbf{p}^{k}(\tau))-\omega_{2}\mathcal{H}(\textbf{p}^{k}(\tau))$ (3)
@@ III-B2 RP Estimation
Assuming that $\mathcal{R}^{k}$ is located in $\textbf{p}^{k}(t)$, a visual feature $\boldsymbol{\phi}(\textbf{p}^{k}(t))$ and a laser scan $\boldsymbol{\psi}(\textbf{p}^{k}(t))$ can be extracted at this location. By comparing $\boldsymbol{\phi}(\textbf{p}^{k}(t))$ and $\boldsymbol{\psi}(\textbf{p}^{k}(t))$ with information stored in neighbors’ FHT-Maps, a single RP estimation can be achieved. The detailed algorithm is presented in Algorithm 1
@@ 4             $r_{\text{match}}$leftarrow $\boldsymbol{\phi}(\textbf{p}^{k}(t))^{T}\boldsymbol{\phi}^{i,j}$
5             if $r_{\text{match}}>th_{\text{match}}$ then
@@ 7                   $T^{k}\leftarrow T^{k}\cup T^{k}_{i}$
In Algorithm 1, $\mathcal{R}^{k}$ compares $\boldsymbol{\phi}(\textbf{p}^{k}(t))$ with visual features in neighbors’ FHT-Map firstly. If a matched main node is found which means $r_{\text{match}}$ is larger than the threshold $th_{\text{match}}$, RP $T^{k}_{i}$ between $\mathcal{R}^{k}$ and $\mathcal{R}^{i}$ can be obtained using global Iterative Closest Point (ICP). In our system, there is no central server and each robot makes decisions locally. Therefore, each robot should have the same global information, for example, estimated RPs, $\mathcal{G}_{r}$, the pose of each robot in its map frame $\mathbf{P}^{k}$, and frontiers of each robot. Since the communication graph $\mathcal{G}_{c}$ is a connected graph, this lightweight information can be broadcast to all robots. So, in Algorithm 1, $T^{k}$ will be shared among all robots by broadcast mechanism.

### v2

@@ I Introduction
Multi-robot systems have been receiving increasing attention in recent years [1, 2]. The cooperative motion in multi-robot systems has been widely studied, for example, consensus [3], exploration [1], and rendezvous [4]. Among the above-mentioned cooperative motion, rendezvous is one of the most important topics, which aims at gathering all robots at a specific location at a given moment [5]. This behavior is common in the real world, for example, a few friends want to meet with each other in a shopping mall or a rescue helicopter is looking for lost hikers in the desert [6].
@@
One of the simplest rendezvous problems involves two robots moving along a straight line [5]. The solutions to this problem can be divided into two types: symmetric and asymmetric. The symmetric version requires the robots to execute the same strategy. Symmetric rendezvous strategy for two robots switching among three fixed locations was proved to be optimal in [13]. Besides, it was extended to a continuous scenario in [7]. The asymmetric version was studied in [14, 15]. Rendezvous for $n>2$ robots on a line was studied in [6] and was extended to 2D scenario in [12, 11]. In these researches, robots rely on stochastic strategies for rendezvous. However, in scenarios with a large number of robots or large-scale environments, the random searching strategy makes rendezvous challenging or even infeasible.
@@
The condition of communication was introduced in [16, 4, 17, 2, 18], where the environment is usually assumed to be known and robots can locate themselves in the environment. Moreover, RPs between robots are available under this assumption. This kind of rendezvous problem has been receiving significant attention in control theory. However, rendezvous in unknown environments is more common and valuable. In real world, the bandwidth is always limited. For instance, satellite-based communication in the wilderness or weak signals in underground parking, their bandwidth can even be less than 1 Mbps. This limitations can lead to time delays in the system, thereby affecting control and decision-making. Therefore, it is important to study rendezvous under the assumption of limited bandwidth communication. As far as we know, this is the first work to perform rendezvous under this assumption.
@@
Recently, multi-robot exploration with unknown initial relative poses (RP) has been studied [19, 1, 20]. In the process, robots determine their RPs and merge the constructed map. In theory, after the exploration and map construction, it is possible to use this full map to achieve rendezvous. However, simply applying multi-robot exploration algorithms to achieve rendezvous is inefficient. Firstly, the exploration aims to construct a complete and detailed map. However, for rendezvous, exploring the environment as minimally as possible is preferred to improve the efficiency, which means an incomplete exploration is needed. Secondly, what kinds of information should be transmitted has also not been studied to effectively achieve RP estimation and rendezvous point selection, which will lead to high communication cost.
@@
Thus, we propose a strategy called partitioned and incomplete exploration for rendezvous (PIER). We focus on how to effectively utilize the limited communication bandwidth (we assume communication range is not limited). This problem involves two aspects specifically: first, how to achieve RP estimation, and second, how to transmit environmental structural information to enable the selection of rendezvous points. In our previous paper, we proposed a lightweight Feature-based Hybrid Topological Map (FHT-Map) [21]. There are two types of nodes: main nodes and support nodes. Main nodes store features extracted from the environment and can facilitate RP estimation. Environmental structural information is stored in support nodes with very low storage requirements, which can be used for rendezvous point selection. We implement the lightweight FHT-Map in this new rendezvous task. By sharing these Maps among robots, we can achieve rendezvous with low bandwidth requirement.
@@
Thus, we divide the rendezvous problem into two steps. Firstly, each robot performs rendezvous based exploration using PIER and constructs its FHT-Map. During this process, robots share the maps among their neighbors and detect similar main nodes for RP estimation, and the FHT-Maps can be merged one by one. Secondly, once the RPs between all robots are determined, the optimal rendezvous point can be selected on the merged map. Our main contributions are:
@@
  • PIER is proposed to perform efficient incomplete exploration with unknown initial RPs.
@@ II Preliminaries
In this paper, we consider the rendezvous problem in a finite 2D space $S\subset\mathbb{R}^{2}$, where the size of the environment is unknown initially. This assumption is easy to satisfy [22], for example, in rendezvous problems for humans within a shopping mall or a campus. We define the k-th robot in the multi-robot system as $\mathcal{R}^{k}$ and the total number of robots is $n$. $\mathbf{P}^{k}=[\mathbf{p}^{k},\mathbf{q}^{k}]\in\mathbf{SE}(2)$ describes the pose of $\mathcal{R}^{k}$ in its own map frame, where $\mathbf{p}^{k}\in\mathbb{R}^{2}$ and $\mathbf{q}^{k}\in\mathbf{SO}(2)$ denote the position and orientation of $\mathcal{R}^{k}$. Besides, a distributed communication framework is used, where all communication are point-to-point and occurs between two robots without a central server. An undirected, unweighted, connected, and non-time-varying communication graph $\mathcal{G}_{c}=(\mathcal{V}_{c},\mathcal{E}_{c})$ is used to describe the communication topology between $n$ robots. The node set is $\mathcal{V}_{c}=\{\mathcal{R}_{1},\cdots,\mathcal{R}_{n}\}$, the edge set is $\mathcal{E}_{c}\subseteq\mathcal{V}_{c}\times\mathcal{V}_{c}$. $\mathcal{N}_{c}^{k}$ defines the set of neighboring robots on $\mathcal{G}_{c}$ for $\mathcal{R}^{k}$.
@@
Different data can be used for RP estimation, for example, images and point clouds. Using raw data from the sensors for RP estimation provides more accurate results but imposes a significant burden on communication. In this paper, we employ lightweight Place Recognition (PR) descriptors [23, 24, 25] for rough RP estimation. Only when two PR descriptors are matched, 2D laser scans of these two points will be transferred for a more accurate RP estimation. A time-varying, undirected and unweighted RP graph $\mathcal{G}_{r}=(\mathcal{V}_{r},\mathcal{E}_{r})$ is used to describe the existence of estimated RPs among $n$ robots. The node set in $\mathcal{G}_{r}$ is $\mathcal{V}_{r}=\{\mathcal{R}^{1},\cdots,\mathcal{R}^{n}\}$, the edge set is $\mathcal{E}_{r}\subseteq\mathcal{V}_{r}\times\mathcal{V}_{r}$. Initially, since there are no estimated RPs between any two robots, $\mathcal{E}_{r}$ is an empty set. When a RP is obtained between robots $m$ and $n$, the edge $(\mathcal{R}^{m},\mathcal{R}^{n})$ is added to $\mathcal{E}_{r}$. Besides, the relative poses between any two robots located on the same subgraph can be obtained through Pose Graph Optimization (PGO), which will be further described in Section III-B.
@@ III PIER Formulation
Robot exploration for a complete reconstruction of the environment has been widely investigated in recent years [20, 28, 1, 29]. In [20], Next Best View (NBV) is used and each robot employs a greedy strategy, enabling rapid acquisition of environmental features but potentially resulting in poor cooperation between robots. In [28, 1], multi-robot exploration is modeled as a multiple traveling salesman problem (m-TSP), which optimizes the movements based on current frontiers. However, m-TSP-based methods are not efficient for the unknown space exploration. Compared with a detailed reconstruction, rendezvous based exploration aims at acquiring environmental features for RP estimation. Therefore, utilizing a greedy strategy to acquire features and considering cooperation between robots will be time-efficient for rendezvous. So, partitioned and incomplete exploration for rendezvous (PIER) is proposed. In PIER, we combine NBV-based and partitioned strategies to achieve rapid cooperative exploration by multiple robots for rendezvous under the condition of unknown initial RPs.
@@ III-A Single Robot NBV Exploration
The NBV-based method employs a greedy strategy, which can achieve a faster coverage for incomplete exploration tasks. During this process, the environment $S$ can be divided into two parts: known space $S_{\text{known}}$ and unknown space $S_{\text{unknown}}$. Furthermore, it can also be classified into two types: free space $S_{\text{free}}$ and occupied space $S_{\text{occ}}$. In PIER, $\mathcal{R}^{k}$ is assigned with a portion of the area $S^{k}$ to explore. The combination of each space forms the environment $S^{1}\cup S^{2}\cdots\cup S^{n}=S$.
@@ III-B RP Estimation between Robots
To achieve rendezvous, each robot first performs NBV-based strategy to explore the unknown environment. During this process, each robot constructs an FHT-Map [21] which can be used for RP estimation. In the FHT-Map, there are two types of nodes: main nodes and support nodes. Main nodes store the visual features $\boldsymbol{\phi}^{k,i}\in\mathbb{R}^{512}$ and geometric features $\boldsymbol{\psi}^{k,i}\in\mathbb{R}^{360}$ of the location, which can be used for RP estimation. When $\mathcal{R}^{k}$ passes a main node of its neighboring robots $\mathcal{N}_{c}^{k}$, it can obtain the RP between $\mathcal{R}^{k}$ and that robot. So, it is crucial to select appropriate positions of main nodes to accelerate RP estimation and rendezvous. In this section, we will describe the details of using FHT-Maps for RP estimation.
@@ III-B1 Selection of Main Nodes
Secondly, since each robot performs incomplete exploration, the probability of each position being visited is different. An intuitive scenario is that robots are more likely to pass through crossroads, which is shown in Fig. 2. Therefore, selecting main nodes at these positions can facilitate RP estimation. We refer to the probability of a location $\textbf{p}^{k}(t)$ being visited as “heat value” $\mathcal{H}(\textbf{p}^{k}(t))\in\mathbb{R}$. We use a neural network to fit the heat value for each location. This network takes the laser scan $\boldsymbol{\psi}(\textbf{p}^{k}(t))$ as input and outputs the fitted $\mathcal{H}(\textbf{p}^{k}(t))$. Heat values of different locations are labeled manually based on Explore-Bench data set [30]. Besides, the backbone of this network is selected as the Pointnet [31].
@@
$\ \boldsymbol{\phi}(\textbf{p}^{k}(\tau))^{T}\boldsymbol{\phi}^{k,i}<\theta_{\text{sim}},\ \forall v_{f}^{k,i}\in\mathcal{V}_{f,\text{main}}^{k},$
where $\theta_{\text{sim}}$ is the similarity threshold between two locations [32]. $\gamma_{1}$ and $\gamma_{2}$ are two parameters which control the number of potential main nodes, which is the same as [21].
In Eq. 2, the feasible set of $\textbf{p}^{k}(\tau)$ is determined firstly. As $\mathcal{R}^{k}$ explores the environment, it gradually moves towards unknown areas, leading to a reduction in main nodes density $\mathcal{D}(\textbf{p}^{k}(t))$. Besides, when $\mathcal{D}(\textbf{p}^{k}(t))$ is smaller than $\gamma_{1}$, the robot will store all features it captures at a fixed rate (30 Hz). These features are called potential main nodes. When $\mathcal{D}(\textbf{p}^{k}(t))$ falls below $\gamma_{2}$, the single main node selection is performed. Firstly, potential main nodes are defined as those who satisfy $\gamma_{2}<\mathcal{D}(\textbf{p}^{k}(\tau))<\gamma_{1}$. Then, we do not expect that the selected main node has high similarity with the existing main nodes in the FHT-Map, which will lead to wrong estimations for RP. Potential main nodes are filtered using $\boldsymbol{\phi}(\textbf{p}^{k}(\tau))^{T}\boldsymbol{\phi}^{k,i}<\theta_{\text{sim}}$. Finally, a new main node will be chosen among the remaining potential main nodes.
Based on Eq. 2, we can obtain a series of Pareto-optimal positions. However, due to that we want to construct a single main node by solving Eq. 2 in a single instance, weighted sum method is used to turn it into a single objective optimization problem, which is called scalarization
$\min_{\tau}\ \omega_{1}\mathcal{D}(\textbf{p}^{k}(\tau))-\omega_{2}\mathcal{H}(\textbf{p}^{k}(\tau)),$ (3)
@@ III-B2 RP Estimation
Assuming that $\mathcal{R}^{k}$ is located in $\textbf{p}^{k}(t)$, a visual feature $\boldsymbol{\phi}(\textbf{p}^{k}(t))$ and a laser scan $\boldsymbol{\psi}(\textbf{p}^{k}(t))$ can be extracted at this location. By comparing $\boldsymbol{\phi}(\textbf{p}^{k}(t))$ and $\boldsymbol{\psi}(\textbf{p}^{k}(t))$ with information stored in neighbors’ FHT-Maps, a single RP estimation can be achieved. The detailed algorithm is presented in Algorithm 1.
@@ 4             $r_{\text{match}}$leftarrow $\boldsymbol{\phi}(\textbf{p}^{k}(t))^{T}\boldsymbol{\phi}^{i,j}$
5             if $r_{\text{match}}>\theta_{\text{match}}$ then
@@ 7                   $T^{k}\leftarrow T^{k}\cup T^{k}_{i}$
In Algorithm 1, $\mathcal{R}^{k}$ compares $\boldsymbol{\phi}(\textbf{p}^{k}(t))$ with visual features in neighbors’ FHT-Map firstly. If a matched main node is found which means $r_{\text{match}}$ is larger than the threshold $\theta_{\text{match}}$, RP $T^{k}_{i}$ between $\mathcal{R}^{k}$ and $\mathcal{R}^{i}$ can be obtained using global Iterative Closest Point (ICP). In our system, there is no central server and each robot makes decisions locally. Therefore, each robot should have the same global information, for example, estimated RPs ($T^{k}$), $\mathcal{G}_{r}$, the pose of each robot in its map frame $\mathbf{P}^{k}$, and frontiers of each robot. Since the communication graph $\mathcal{G}_{c}$ is a connected graph, these lightweight information can be broadcast to all robots.